OpenStudy (anonymous):

Find the GCF of 30x^3, 40x^5

OpenStudy (anonymous):

\[30x^3 = 2 \times 3 \times 5 \times x \times x \times x\] \[40x^5 = 2 \times 2 \times 2 \times 5 \times x \times x \times x \times x \times x\] common factors of 30x^3 and 40x^5 is 2×5×x×x×x therefore GCF of 30x^3, 40x^5 = 2×5×x×x×x = 10x^3 @Katelynmorris

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

@Katelynmorris medal please